# Attitude Trajectories Stabilizable by Linear Control

### v1

@@ III Linear constant difference attitude control
IV Validation using Simulations
@@
In this section, we present simulation results to validate the key ideas presented in the previous section. The stability of the system is characterized by the eigenvalues of the state matrix in (22), which are the roots of the characteristic polynomial in (28). As we saw in the previous section, the best we can hope with respect to stability is that all the eigenvalues lie on the imaginary axis (since the trace of (22) is zero). In order that the $\lambda$ of (28) lie on the imaginary axis, the $\lambda^{\prime}$ of (29) must be negative real numbers. The latter condition may be verified by evaluating the discriminant (for realness) and applying the Routh-Hurwitz test (for negativity).
@@
The first result is to demonstrate that all roots of (29) are in the left-half of the complex plane. This can be easily accomplished using the Routh-Hurwitz test (see Routh-Hurwitz stability criterion in, for example, [19]). Since $|e_{0}|\leq 1$, it is obvious that the coefficients of (29) are all non-negative. We also need to ensure that the product of the coefficients $a$ and $b$ of $\lambda^{\prime 2}$ and $\lambda^{\prime}$ is greater than the constant coefficient $c$. This condition is verified in Fig. 1.
Figure 1: The product of the coefficients $a$ and $b$ of $\lambda^{\prime 2}$ and $\lambda^{\prime}$ in (29) must be greater than the constant coefficient $c$ in order to ensure all roots lie in the left half of the complex plane.
The next plot in Fig. 2 shows that the discriminant $\Delta$ of (29) changes its sign from positive to negative when $e_{0}\approx 0.84$. This marks the boundary of the unstable and marginally stable regions with respect to the error angle.
Figure 2: The discriminant of (29) may be plotted to graphically determine the critical error angle that separates the marginally stable region from the unstable region for the linearized dynamics in (21). Here, we see that the plot for the discriminant crosses the $x$-axis at $e_{0}$ approximately equal to 0.85, which prognosticates instability for values of $e_{0}$ less than approximately 0.84.
We now verify the cumulative outcome of the previous two plots, namely the condition that the eigenvalues $\lambda$ of (28) are in the left-half of the complex plane. Fig. 3 shows the locus of each of the six eigenvalues of the system matrix $A$ in (22). We can see that all six eigenvalues are indeed purely imaginary until $e_{0}$ falls below approximately 0.85. For $e_{0}$ less than approximately 0.84, the system has two purely imaginary eigenvalues, and four complex eigenvalues. Of the latter four, two lie in the positive half of the complex plane and hence indicate instability.
Figure 3: The root locus for the system matrix in (22) shows the movement of all six eigenvalues of the state matrix in. We can see that the linearized system whose dynamics are governed by (21) is marginally stable when $e_{0}\gtrsim 0.85$.
V Linear Attitude Control Design
We will now design a linear attitude controller using the marginally stable natural dynamics in (21) as a starting point, and by including feedback terms from the measured or estimated attitude and angular velocity errors, to move the eigenvalues of the system into the left half of the complex plane. To this end, we include four gain parameters, $K_{1}$ to $K_{4}$, in the feedback law as follows:
$\begin{bmatrix}\dot{\tilde{e}}_{v}\\ \dot{\tilde{\omega}}\end{bmatrix}$ $=A_{fb}\begin{bmatrix}\tilde{e}_{v}\\ \tilde{\omega}\end{bmatrix}$
where the feedback matrix $A_{fb}$ may be expressed in terms of the matrix $A$ in (22) as
$A+\begin{bmatrix}0&0\\ K_{1}1_{3\times 3}+K_{2}[w\times]&K_{3}1_{3\times 3}+K_{4}[e_{v}\times]\end{bmatrix}.$ (31)
The gains $K_{1}$ and $K_{3}$ provide the usual proportional and derivative feedback from the errors, while the gains $K_{2}$ and $K_{4}$ are introduced to take care of the cross-product terms in the error dynamics described in (20). Since we are considering linear control, $e_{v}$ and $w$ in (31) would be the nominal attitude error and the nominal body-frame (half) angular velocity.
Figure 4: The eigenvalues of the linearized attitude controller may all be moved into the left half of the complex plane by actively using negative feedback. The plots above correspond to a quaternion error of $0.11-e_{0}=$ 0.1, and use tuned feedback gains of $K_{1}=K_{3}=-0.3$, $K_{2}=1$, and $K_{4}=-0.5$.
The gains in (31) may now be tuned using simulations to stabilize the system by locating the eigenvalues in the left half of the complex plane resulting in the following values: $K_{1}=K_{3}=-0.3$, $K_{2}=1$, and $K_{4}=-0.5$. The root locus of the system, for feedback gains scaled by a factor varying from a tenth to ten times, is shown in Fig. 4.

### v2

@@ III-B Linearized (Perturbation) Dynamics
IV Attitude Control Design
IV-A Nominal Trajectory Parameterization
Without loss of generality, a nominal constant quaternion difference trajectory for the desired and body attitudes $\check{p}$ and $\check{q}$ may be parameterized (modulo a constant rotation isomorphism) in terms of two scalar signals $\alpha(t)$ and $\beta(t)$, and a scalar constant $\phi$ as follows:
$\check{p}$ $=\begin{bmatrix}c_{\phi}c_{\alpha}\\ c_{\phi}s_{\alpha}c_{\beta}\\ c_{\phi}s_{\alpha}s_{\beta}\\ -s_{\phi}\end{bmatrix},\;\check{q}=\begin{bmatrix}c_{\phi}c_{\alpha}\\ c_{\phi}s_{\alpha}c_{\beta}\\ c_{\phi}s_{\alpha}s_{\beta}\\ s_{\phi}\end{bmatrix}$, (34)
$\Rightarrow\check{e}$ $=\check{p}^{-1}\otimes\check{q}$
$=\begin{bmatrix}c_{2\phi}\\ -s_{2\phi}s_{\alpha}s_{\beta}\\ s_{2\phi}s_{\alpha}c_{\beta}\\ s_{2\phi}c_{\alpha}\end{bmatrix}=c_{2\phi}\check{1}+s_{2\phi}\begin{bmatrix}0\\ -s_{\alpha}s_{\beta}\\ s_{\alpha}c_{\beta}\\ c_{\alpha}\end{bmatrix}$. (35)
Let us assume that $\alpha(t)$ and $\beta(t)$ are arbitrary twice-differentiable functions of time. Taking the derivatives, we obtain the desired and body (half-)angular velocities $v$ and $w$ in terms of the constant parameter $\phi$, and the signals $\alpha(t)$ and $\beta(t)$:
$\dot{\check{p}}$ $=\dot{\check{q}}=\dot{\alpha}c_{\phi}\begin{bmatrix}-s_{\alpha}\\ c_{\alpha}c_{\beta}\\ c_{\alpha}s_{\beta}\\ 0\end{bmatrix}+\dot{\beta}c_{\phi}s_{\alpha}\begin{bmatrix}0\\ -s_{\beta}\\ c_{\beta}\\ 0\end{bmatrix}$ (36)
$\vec{v\mkern-1.0mu}$ $=\check{p}^{-1}\otimes\dot{\check{p}}$
$=(\dot{\alpha}c_{\phi}^{2}-\dot{\beta}c_{\phi}s_{\phi}s_{\alpha})\begin{bmatrix}c_{\beta}\\ s_{\beta}\\ 0\end{bmatrix}+(\dot{\alpha}c_{\phi}s_{\phi}+\dot{\beta}c_{\phi}^{2}s_{\alpha})\begin{bmatrix}-c_{\alpha}s_{\beta}\\ c_{\alpha}c_{\beta}\\ -s_{\alpha}\end{bmatrix}$. (37)
Since $\check{q}$ and $\vec{w\mkern-1.0mu}$ are obtained from $\check{p}$ and $\vec{v\mkern-1.0mu}$ by replacing $\phi$ with its negative $\phi$, we must have
$\vec{w\mkern-1.0mu}$ $=(\dot{\alpha}c_{\phi}^{2}+\dot{\beta}c_{\phi}s_{\phi}s_{\alpha})\begin{bmatrix}c_{\beta}\\ s_{\beta}\\ 0\end{bmatrix}+(\dot{\beta}c_{\phi}^{2}s_{\alpha}-\dot{\alpha}c_{\phi}s_{\phi})\begin{bmatrix}-c_{\alpha}s_{\beta}\\ c_{\alpha}c_{\beta}\\ -s_{\alpha}\end{bmatrix}$. (38)
We can see from (35), (37) and (38) that $\vec{e\mkern-1.0mu}_{\text{v}}^{\text{T}}\vec{v\mkern-1.0mu}=\vec{e\mkern-1.0mu}_{\text{v}}^{\text{T}}\vec{w\mkern-1.0mu}=0$, which motivates us to define the orthonormal triad
$\begin{bmatrix}\hat{x}&\hat{y}&\hat{z}\end{bmatrix}$ $=\begin{bmatrix}c_{\beta}&-c_{\alpha}s_{\beta}&-s_{\alpha}s_{\beta}\\ s_{\beta}&c_{\alpha}c_{\beta}&s_{\alpha}c_{\beta}\\ 0&-s_{\alpha}&c_{\alpha}\end{bmatrix}$
$=\begin{bmatrix}c_{\beta}&-s_{\beta}\\ s_{\beta}&c_{\beta}\\ &&1\end{bmatrix}\begin{bmatrix}1\\ &c_{\alpha}&s_{\alpha}\\ &-s_{\alpha}&c_{\alpha}\end{bmatrix}$, (39)
with derivatives
[…]
$\dot{\hat{y}}$ $=-\dot{\alpha}\hat{z}-\dot{\beta}c_{\alpha}\hat{x}$, (41)
$\dot{\hat{z}}$ $=-\dot{\beta}s_{\alpha}\hat{x}+\dot{\alpha}\hat{y}$, (42)
$\vec{\omega\mkern-1.0mu}$ $=-\dot{\alpha}\hat{x}-\dot{\beta}s_{\alpha}\hat{y}+\dot{\beta}c_{\alpha}\hat{z}$. (43)
From (37) and (38), the magnitudes of the nominal $\vec{v\mkern-1.0mu}$ and $\vec{w\mkern-1.0mu}$ are
$\vec{v\mkern-1.0mu}^{\text{T}}\vec{v\mkern-1.0mu}$ $=\vec{w\mkern-1.0mu}^{\text{T}}\vec{w\mkern-1.0mu}=(\dot{\alpha}c_{\phi}^{2}\mp\dot{\beta}c_{\phi}s_{\phi}s_{\alpha})^{2}+(\dot{\beta}c_{\phi}^{2}s_{\alpha}\pm\dot{\alpha}c_{\phi}s_{\phi})^{2}$
$=\dot{\alpha}^{2}c_{\phi}^{2}+\dot{\beta}^{2}c_{\phi}^{2}s_{\alpha}^{2}.$ (44)
The desired (half-) angular acceleration $\dot{\vec{v\mkern-1.0mu}}$ may also be expressed in terms of the parameter $\phi$ and the signals $\alpha(t)$ and $\beta(t)$ as
$\dot{\vec{v\mkern-1.0mu}}$ $=(\ddot{\alpha}c_{\phi}^{2}-2\dot{\alpha}\dot{\beta}c_{\phi}s_{\phi}c_{\alpha}-\ddot{\beta}c_{\phi}s_{\phi}s_{\alpha}-\dot{\beta}^{2}c_{\phi}^{2}c_{\alpha}s_{\alpha})\hat{x}$
$\;+(\ddot{\alpha}c_{\phi}s_{\phi}+2\dot{\alpha}\dot{\beta}c_{\phi}^{2}c_{\alpha}+\ddot{\beta}c_{\phi}^{2}s_{\alpha}-\dot{\beta}^{2}c_{\phi}s_{\phi}c_{\alpha}s_{\alpha})\hat{y}$
$\;-(\dot{\alpha}^{2}+\dot{\beta}^{2}s_{\alpha}^{2})c_{\phi}s_{\phi}\hat{z}$. (45)
The body (half-) angular acceleration $\dot{\vec{w\mkern-1.0mu}}$ may be obtained by substituting $\phi$ with $\phi$ in (45).
IV-B Feedback and Control of General Attitude Trajectories
For general desired attitude trajectories that do not satisfy the requirement in (15) for constant quaternion difference dynamics, we cannot directly use feedforward parameterization and control of the body attitude presented in the previous subsection. However, the conditions in (13) and (16) can still be used advantageously to design an efficient attitude controller as described in this subsection.
Since equations (13) and (16) together impose four scalar degrees of constraint on the three scalar coordinates of $w$, we could use feedback to define two independent control laws for $\dot{\vec{w\mkern-1.0mu}}$
$\vec{e\mkern-1.0mu}_{\text{v}}^{\text{T}}\dot{\vec{w\mkern-1.0mu}}$ $=g(\vec{v\mkern-1.0mu},\dot{\vec{v\mkern-1.0mu}},\check{e},\vec{w\mkern-1.0mu})$
$=-\vec{w\mkern-1.0mu}^{\text{T}}\dot{\vec{e\mkern-1.0mu}}_{\text{v}}-k\vec{e\mkern-1.0mu}_{\text{v}}^{\text{T}}\vec{w\mkern-1.0mu}$
$=e_{0}\vec{w\mkern-1.0mu}^{\text{T}}(\vec{v\mkern-1.0mu}-\vec{w\mkern-1.0mu})-\vec{e\mkern-1.0mu}_{\text{v}}^{\text{T}}\vec{v\mkern-1.0mu}\times\vec{w\mkern-1.0mu}-k\vec{e\mkern-1.0mu}_{\text{v}}^{\text{T}}\vec{w\mkern-1.0mu},$ (46)
$\dot{\vec{w\mkern-1.0mu}}$ $=\vec{h\mkern-1.0mu}(\vec{v\mkern-1.0mu},\dot{\vec{v\mkern-1.0mu}},\check{e},\vec{w\mkern-1.0mu})$
$=\dot{e}_{0}\vec{v\mkern-1.0mu}+e_{0}\dot{\vec{v\mkern-1.0mu}}-\dot{\vec{e\mkern-1.0mu}}_{\text{v}}\times\vec{v\mkern-1.0mu}-\vec{e\mkern-1.0mu}_{\text{v}}\times\dot{\vec{v\mkern-1.0mu}}-L(\vec{w\mkern-1.0mu}-e_{0}\vec{v\mkern-1.0mu}+\vec{e\mkern-1.0mu}_{\text{v}}\times\vec{v\mkern-1.0mu})$
$=e_{0}\dot{\vec{v\mkern-1.0mu}}-\vec{e\mkern-1.0mu}_{\text{v}}\times\dot{\vec{v\mkern-1.0mu}}-\vec{v\mkern-1.0mu}\,\vec{e\mkern-1.0mu}_{\text{v}}^{\text{T}}\vec{w\mkern-1.0mu}-\vec{w\mkern-1.0mu}\,\vec{e\mkern-1.0mu}_{\text{v}}^{\text{T}}\vec{v\mkern-1.0mu}+e_{0}\vec{v\mkern-1.0mu}\times\vec{w\mkern-1.0mu}+\vec{e\mkern-1.0mu}_{\text{v}}\vec{v\mkern-1.0mu}^{\text{T}}(\vec{v\mkern-1.0mu}+\vec{w\mkern-1.0mu})$
$\quad-L(\vec{w\mkern-1.0mu}-e_{0}\vec{v\mkern-1.0mu}+\vec{e\mkern-1.0mu}_{\text{v}}\times\vec{v\mkern-1.0mu}).$ (47)
where $k\in\mathbb{R}$ is a positive feedback gain and $L\in\mathbb{R}^{3\times 3}$ is a positive definite feedback matrix. We may design the overall control input as follows:
$\dot{\vec{w\mkern-1.0mu}}^{\star}$ $=\vec{h\mkern-1.0mu}+(1-r)\frac{(g-\vec{e\mkern-1.0mu}_{\text{v}}^{\text{T}}\vec{h\mkern-1.0mu})\vec{e\mkern-1.0mu}_{\text{v}}}{\vec{e\mkern-1.0mu}_{\text{v}}^{\text{T}}\vec{e\mkern-1.0mu}_{\text{v}}},$ (48)
where $r$ is a tunable gain parameter. When $r=0$, we identically have $\vec{e\mkern-1.0mu}_{\text{v}}^{\text{T}}\dot{\vec{w\mkern-1.0mu}}^{\ast}=0$, and when $r=1$, we identically have $\dot{\vec{w\mkern-1.0mu}}^{\ast}=\vec{h\mkern-1.0mu}$. One remaining flaw with the control law in (48) is that $g$ (defined in (46)) could remain nonzero as $\check{e}$ approaches identity and $\vec{e\mkern-1.0mu}_{\text{v}}$ approaches zero, leading to unbounded control for $\dot{\vec{w\mkern-1.0mu}}$ in (48). A simple remedy is to replace $\vec{w\mkern-1.0mu}$ with $e_{0}\vec{v\mkern-1.0mu}-\vec{e\mkern-1.0mu}_{\text{v}}\times\vec{v\mkern-1.0mu}$ in $\dot{\vec{e\mkern-1.0mu}}_{\text{v}}$ in $g$ as follows
$g$ $=e_{0}\vec{w\mkern-1.0mu}^{\text{T}}((1-e_{0})\vec{v\mkern-1.0mu}+\vec{e\mkern-1.0mu}_{\text{v}}\times\vec{v\mkern-1.0mu})-\vec{e\mkern-1.0mu}_{\text{v}}^{\text{T}}\vec{v\mkern-1.0mu}\times\vec{w\mkern-1.0mu}-k\vec{e\mkern-1.0mu}_{\text{v}}^{\text{T}}\vec{w\mkern-1.0mu},$ (49)
yielding a bounded control law even when $\check{e}\rightarrow\check{1}$.
Rather than take the weighted average of two acceleration control laws as done above, a theoretically more fruitful approach yielding a strong stability result is to first weigh the constraints on the velocities, and then differentiate the combination to derive the acceleration control law as done below.
Let the target angular velocity be
$\vec{w\mkern-1.0mu}^{\ast}$ $=e_{0}\vec{v\mkern-1.0mu}-\vec{e\mkern-1.0mu}_{\text{v}}\times\vec{v\mkern-1.0mu}-(1-r)\frac{e_{0}\vec{e\mkern-1.0mu}_{\text{v}}\vec{e\mkern-1.0mu}_{\text{v}}^{\text{T}}\vec{v\mkern-1.0mu}}{|\vec{e\mkern-1.0mu}_{\text{v}}|}$ (50)
where $r$ is a weighting factor between the constraints in (13) and (16). However, (13) is only approximately satisfied for $|\vec{e\mkern-1.0mu}_{\text{v}}|\neq 1$. Equation (13) could have been identically satisfied by using the combination
$\vec{w\mkern-1.0mu}^{\ast}$ $=e_{0}\vec{v\mkern-1.0mu}-\vec{e\mkern-1.0mu}_{\text{v}}\times\vec{v\mkern-1.0mu}-(1-r)\frac{e_{0}\vec{e\mkern-1.0mu}_{\text{v}}\vec{e\mkern-1.0mu}_{\text{v}}^{\text{T}}\vec{v\mkern-1.0mu}}{|\vec{e\mkern-1.0mu}_{\text{v}}|^{2}}$ (51)
with $|\vec{e\mkern-1.0mu}_{\text{v}}|^{2}$ replacing the $|\vec{e\mkern-1.0mu}_{\text{v}}|$ in the denominator of the third term on the right hand side of (50), but this new law has the even bigger problem of diverging when $|\vec{e\mkern-1.0mu}_{\text{v}}|\rightarrow 0$ upon differentiation, which is an important consideration for nonlinear control design.
Returning to (50), we see that
$\lim_{r\rightarrow 0}\vec{e\mkern-1.0mu}_{\text{v}}^{\text{T}}\vec{w\mkern-1.0mu}^{\ast}$ $=e_{0}(1-|\vec{e\mkern-1.0mu}_{\text{v}}|)\vec{e\mkern-1.0mu}_{\text{v}}^{\text{T}}\vec{v\mkern-1.0mu},$
$\lim_{r\rightarrow 1}\vec{w\mkern-1.0mu}^{\ast}$ $=e_{0}\vec{v\mkern-1.0mu}-\vec{e\mkern-1.0mu}_{\text{v}}\times\vec{v\mkern-1.0mu},$
which satisfy (13) approximately and (16) exactly. The residual noncompliance of (13) is of the same order of magnitude as the residual noncompliance of (15) in the desired trajectory. The acceleration control law from (50) is therefore
$\dot{w}$ $=\dot{e}_{0}\vec{v\mkern-1.0mu}+e_{0}\dot{\vec{v\mkern-1.0mu}}-\dot{\vec{e\mkern-1.0mu}}_{\text{v}}\times\vec{v\mkern-1.0mu}-\vec{e\mkern-1.0mu}_{\text{v}}\times\dot{\vec{v\mkern-1.0mu}}-(1-r)\frac{\vec{e\mkern-1.0mu}_{\text{v}}\vec{e\mkern-1.0mu}_{\text{v}}^{\text{T}}}{|\vec{e\mkern-1.0mu}_{\text{v}}|}(\dot{e}_{0}\vec{v\mkern-1.0mu}+e_{0}\dot{\vec{v\mkern-1.0mu}})$
$\quad-(1-r)\frac{e_{0}(\dot{\vec{e\mkern-1.0mu}}_{\text{v}}\vec{e\mkern-1.0mu}_{\text{v}}^{\text{T}}+\vec{e\mkern-1.0mu}_{\text{v}}\dot{\vec{e\mkern-1.0mu}}_{\text{v}}^{\text{T}})\vec{v\mkern-1.0mu}}{|\vec{e\mkern-1.0mu}_{\text{v}}|}+(1-r)\frac{e_{0}(\vec{e\mkern-1.0mu}_{\text{v}}^{\text{T}}\dot{\vec{e\mkern-1.0mu}}_{\text{v}})\vec{e\mkern-1.0mu}_{\text{v}}\vec{e\mkern-1.0mu}_{\text{v}}^{\text{T}}\vec{v\mkern-1.0mu}}{|\vec{e\mkern-1.0mu}_{\text{v}}|^{3}}$
$\quad-L\left[\vec{w\mkern-1.0mu}-e_{0}\vec{v\mkern-1.0mu}+\vec{e\mkern-1.0mu}_{\text{v}}\times\vec{v\mkern-1.0mu}+(1-r)\frac{e_{0}\vec{e\mkern-1.0mu}_{\text{v}}\vec{e\mkern-1.0mu}_{\text{v}}^{\text{T}}\vec{v\mkern-1.0mu}}{|\vec{e\mkern-1.0mu}_{\text{v}}|}\right].$ (52)
In the above control law, $\dot{e}_{0}$ and $\dot{\vec{e\mkern-1.0mu}}_{\text{v}}$ are given in terms of $e_{0}$, $\vec{e\mkern-1.0mu}_{\text{v}}$, $\vec{v\mkern-1.0mu}$, and $\vec{w\mkern-1.0mu}$ through (18).
Calling the right hand side of (50) as $\vec{f\mkern-1.0mu}$, we have
$\dot{\vec{w\mkern-1.0mu}}$ $=\dot{\vec{f\mkern-1.0mu}}-L(\vec{w\mkern-1.0mu}-\vec{f\mkern-1.0mu}).$
Defining a positive definite candidate Lyapunov function
$V_{L}$ $=\frac{1}{2}|\vec{w\mkern-1.0mu}-\vec{f\mkern-1.0mu}|^{2},$ (53)
we obtain for its first derivative
$\dot{V}_{L}$ $=-(\vec{w\mkern-1.0mu}-\vec{f\mkern-1.0mu})^{\text{T}}L(\vec{w\mkern-1.0mu}-\vec{f\mkern-1.0mu}),$ (54)
which is a negative definite function with bounded second derivative $\ddot{V}_{L}=-(\vec{w\mkern-1.0mu}-\vec{f\mkern-1.0mu})^{\text{T}}L()$, which implies asymptotic convergence of $\vec{w\mkern-1.0mu}$ to $\vec{f\mkern-1.0mu}$ using Barbalat’s Lemma. It should now be clear why we insisted on choosing the bounded control law of (52).
V Validation using Simulations
@@
In this section, we present simulation results to validate the key ideas presented in the previous section. The stability of the constant quaternion difference system is characterized by the eigenvalues of the state matrix in (24), which are the roots of the characteristic polynomial in (31). As we saw in the previous section, the best we can hope with respect to stability is that all the eigenvalues lie on the imaginary axis (since the trace of (24) is zero). In order that the $\lambda$ of (31) lie on the imaginary axis, the $\lambda^{\prime}$ of (32) must be negative real numbers. The latter condition may be verified by evaluating the discriminant (for realness) and applying the Routh-Hurwitz test (for negativity).
@@
The first result is to demonstrate that all roots of (32) are in the left-half of the complex plane for $e_{0}>0$. This can be easily accomplished using the Routh-Hurwitz test (see Routh-Hurwitz stability criterion in, for example, [Dorf:11a]). For $0<e_{0}\leq 1$, it is obvious that the coefficients of (32) are all non-negative. We also need to ensure that the product of the coefficients $a$ and $b$ of $\lambda^{\prime 2}$ and $\lambda^{\prime}$ is greater than the constant coefficient $c$. This condition is verified on the left side of Fig. 1. The adjacent plot on the right side shows that the discriminant $\Delta$ of (32) also remains positive for all $e_{0}>0$. Together, the plots show that we obtain marginal stability for $e_{0}>0$. For $e_{0}\leq 0$, the system is obviously unstable.
Figure 1: Left: The product of the coefficients $a$ and $b$ of $\lambda^{\prime 2}$ and $\lambda^{\prime}$ in (32) must be greater than the constant coefficient $c$ in order to ensure all roots lie in the left half of the complex plane. Right: The discriminant of (32) may be plotted to graphically determine the critical error angle that separates the marginally stable region from the unstable region for the linearized dynamics in (23). Here, we see that both plots indicate marginal stability of the constant quaternion difference system for $e_{0}>0$ (error angle less than $\pi$).
We next verify the control design in Section IV based upon the stability analysis in Section III. For the prescribed attitude, we use the nominal trajectory presented in Subsection IV-A. For the actual trajectory, we begin with an initial error of one radian, which is less than $\pi$, so $e_{0}>0$ as required by the stability analysis. The controller gain matrix $L$ is chosen as the identity matrix while the weight $r$ is chosen as 0.5. The error in the attitude and the violation in the constraints (13) and (16) are plotted in the figure below. We can clearly see asymptotic convergence of the error to zero.
Figure 2: Left: The quaternion difference error can be seen to asymptotically converge to zero with the control law proposed in (52). Right: The constraint violations also approach zero.